### v1

@@ 4.2 Compute-I/O Overlapping
LMCache’s data movement often happens in parallel to LLM inference computation. For example, during decoding, a KV cache is computed while being stored to storage devices at the same time. Naive data movement operations can block the LLM inference computations, or vice versa. LMCache employs multiple optimizations aiming for overlapping LLM inference computations with I/O.
@@
Asynchronous compute & prefetch:  In many scenarios, there is a time gap between when the inference scheduler admits a query and when the query’s KV cache is actually needed for inference. For example, if 100 queries arrive simultaneously but the inference engine can only process 50 in parallel, the remaining 50 must wait in the queue before their inference begins. LMCache exploits this idle interval to prefetch the queued queries’ KV cache from slower storage tiers into faster ones (e.g., from remote disk to local CPU memory). As a result, when the actual inference computation starts, the required KV cache can be loaded directly from faster-tier storage, significantly reducing loading delay. LMCache allows users to configure the target tier for prefetching based on their own needs in latency SLO and resource constraints.
Process separation:  Co-locating LMCache ’s data movement with the inference engine in a single process introduces 5%–10% latency overhead due to CPU resource contention. To eliminate this overhead, LMCache decouples data movement into a separate process, isolating it from the inference computation process. Additionally, process separation also enables memory sharing across multiple inference engine instances. When data movement is co-located with computation, each inference instance maintains its own CPU memory pool, so instances cannot access one another’s KV cache without inter-process communications. By contrast, a standalone data movement process manages a unified CPU memory pool, allowing multiple inference instances to store and load KV cache from the same pool, thereby improving efficiency and reducing redundancy.
@@ 4.3 Minimum Data Copy
Zero-Copy Operations:  When transferring KV cache to remote storage or another GPU instance (e.g., in disaggregated prefill), LMCache minimizes data duplication through a reference counter. Specifically, when KV cache is written to multiple destinations—such as local CPU memory, local disk, or remote disk—LMCache increments a reference counter for the shared data instead of creating new copies. Each completed write decrements the counter, and once the count reaches zero, the data is released. This design ensures that data is shared across concurrent write operations without unnecessary replication, thus reducing memory pressure and improving efficiency. This technique is simlar to PCB counter in operating systems Strecker (1978).
@@
get_num_new_matched_tokens(query) $\rightarrow$ matched_tokens Returns the number of cache-hit tokens found in LMCache ’s backend.
@@
Note that if a query attempts to allocate pages beyond the current pointer, the allocation must stall until the current pointer moves right enough to cover the required pages. Thus, a key trade-off in this design is that: the number of duplicated pages—i.e., the region defined by end pointerstart pointer—between GPU and CPU memory. A smaller duplication window reduces the duplication ratio but increases the likelihood of allocation stalls. For instance, if only one page is duplicated and an inference query requires three pages, the query must wait until the current and end pointers advance by two additional pages. On the other hand, if three pages are duplicated, the same query can proceed immediately without stalling, though at the cost of higher duplication ratio.
@@ 5 Standardized Interface for Connecting the KV Caching Layer and Inference Engine
To address this challenge, LMCache introduces a standardized KV cache connector interface that decouples KV cache management from the inference engine backend. This design ensures that LMCache remains compatible regardless of how the underlying inference engine evolves.
Concretely, in vLLM, LMCache integrates through two key interfaces: 1) the scheduler, where the number of prefill tokens directly influences scheduling decisions and are changed by LMCache (i.e., if there is cache hit in LMCache, the number of tokens that need to be newly prefilled changes); and 2) the model runner, where inference computation occurs, and where KV cache loading and storing must be performed before and after execution.
@@
Putting it together, when a query comes in, the scheduler first calls get_num_new_matched_tokens which queries LMCache to see cache hit tokens in the backend. Then the update_state_after_alloc function decides whether each page in vLLM needs to be loaded from external storage backend based on matched tokens information from LMCache. If the cache hit tokens are greater than zero, build_connector_meta function is called to prepare necessary metadata to load or store KV cache from storage devices.
@@
Interface Description
lookup(tokens) $\rightarrow$ {instance_id: hit_tokens} Returns the instance names of KV caches containing prefix-match tokens with the given token list.
query_ip(instance_ids) $\rightarrow$ IP Returns the IP addresses for the corresponding instance_ids
move(source, destination, tokens) Moves the KV cache of the given token list from a source instance to destination.
clear(tokens, instance_id, storage_device) Clears the KV cache for corresponding tokens from the storage_device in instance.
pin(tokens, instance, storage_device) Pins the KV cache for corresponding tokens at the storage_device in instance.
compress(tokens, instance, storage_device, compression_method) Compresses the KV cache for the corresponding tokens with a compression method and store it at the storage_device in instance.
Table 3: Functions in LMCache’s connector.
@@ 6 Controller Interfaces
There are many scenarios where higher-level applications need to see KV cache locations and operate on that. For example, query routing can benefit from knowing where the largest number of prefix cache hit tokens is at. To support such use cases, LMCache exposes APIs, as shown in Table 3, for querying and manipulating KV cache, such as locating cached entries or migrating them across instances.
The KV cache controller in LMCache consists of two components: a centralized manager and per-instance workers. Incoming controller queries are handled by the centralized manager, which dispatches the appropriate operations to workers running within each LMCache instance. The remainder of this section illustrates how to use the controller interfaces through several example applications.
KV cache–aware routing:  In this case, higher-level query routers would like to direct queries to instances with the largest number of prefix cache hits. The router first calls lookup(tokens), where the centralized manager dispatches the lookup queries to individual LMCache instances, to identify the instances containing matching prefix tokens, then invokes query_ip(instance_ids) to map instance ids to IP addresses. Finally, the query is forwarded to the IP with the highest number of hits.
KV cache migration:  When an instance holding KV cache fails or load balancing is required, the KV cache may need to be migrated across instances. The move(source, destination, tokens) interface moves the KV cache corresponding to tokens from the source instance to the destination.
@@
KV cache clearance:  Applications may clear cache when switching models or reclaiming memory. The clear(tokens, instance, location) function removes the KV cache corresponding to tokens stored at a specific storage location (e.g., GPU memory or CPU memory) within the given instance.
Pinning KV cache in GPU memory:  Some contexts, such as system prompts in chatbot applications, should remain in GPU memory for fast access. The pin(instance, location, tokens) function pins the specified KV cache in the chosen storage tier (GPU memory) of the given instance.
@@
There are other more advanced interfaces users can call to manipulate the KV cache, such as the compress(tokens, instance, location, compression_method) function which compresses KV cache stored in a specific location in a LMCache instance with compression_method so that the KV cache takes smaller amount of space in storage. Users can freely call these interfaces to explicitly manage the KV cache in their own applications based on their needs.
@@ 7 Evaluation
Figure 5: Single-node evaluation results.
Figure 6: Real-trace evaluation results.
Figure 7: Remote backend evaluation results.
Figure 8: TTFT and ITL CDF comparison for PD.
Figure 9: Comparing to SGLang CPU Offloading result.
Figure 11: Latency breakdown of PD.
@@ 7.1 Setup
Scenario Acronym
Single-node /
Multi-node
Network
Medium
Real-world
Examples
CPU Offload Single-node N.A.
Single-node
CPU Offloading
Central Storage Single-node Ethernet
Centralized
Storage Server
PD Single-node NVLink
PD
Disaggregation
@@
Models:  We compare LMCache against baseline solutions on popular open source models adopted by industry: meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-Coder-32B-Instruct, Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8, Qwen/Qwen2.5-72B-Instruct.
Datasets:  LMCache is evaluated on several datasets, including emulated multi-round question answering, long context question answering from LongBench Bai et al. (2024), and random dataset from vLLM official benchmarking script Kwon et al. (2023c).
@@
Baselines:  We compare LMCache with several baselines, including 1) vLLM with prefix caching, which implements prefix caching, but only keeps limited amount of KV cache inside GPU memory; 2) commercial offering A, B, and C, which provide dedicated endpoint service that reserves GPUs for users to run a user-defined model.
@@ 7.2 Single-node CPU Offloading
Understanding LMCache’s gains:  LMCache outperforms baselines for several reasons. Compared with naïve vLLM prefill, which caches KV data only in GPU memory, LMCache leverages CPU offloading. Since CPU memory can hold far more KV cache than GPU memory, LMCache achieves significantly higher cache hit ratios. With its efficient CPU-to-GPU loading scheme, cached data can be fetched quickly to accelerate inference. Our comparison with closed-source commercial alternatives is conducted in a black-box manner since their internal implementations are not publicly available. From the end-to-end results, we hypothesize that Commercial Option #1 lacks a KV cache offloading mechanism to secondary storage. In contrast, Commercial Option #2 likely supports KV cache offloading to secondary storage, yet its performance is still worse than LMCache.
@@ 7.3 Centralized Storage Server
Next, we run LMCache for KV cache sharing through a centralized remote server, that is connected to the GPU instance with a bandwidth of 15 Gbps, following the setup of central storage in Table 4. For this experiment, we evaluate using the TriviaQA dataset from LongBench Bai et al. (2024), a widely adopted benchmark for long-context evaluation. We follow the official vLLM benchmarking scripts Kwon et al. (2023c), which generate inference queries according to a Poisson distribution at a specified QPS.
@@
We note, however, that loading KV cache from the remote backend introduces greater latency than loading from CPU memory, since the remote backend has a much lower bandwidth. As a result, the loading delay may even surpass the prefill delay, particularly when the input context is short or model is small, as prefilling is too fast in such cases—a scenario that we will demonstrate later in 7.7.
@@ 7.4 PD disaggregation
7.5 Real-trace Driven Evaluation
We also evaluate LMCache on a real production trace provided by company FF [^4] Anonymized due to NDA.. The trace includes query arrival timestamps and actual multi-round chat inputs. Notably, later rounds within the same user session often reuse conversation history from earlier rounds, resulting in prefix reuse. On average, the inputs in the trace are about 4K tokens long.
Since we do not have access to company $F$’s proprietary models, we replay the trace using the Sao10K/L3-8B-Lunaris-v1 model. To make the experiment tractable, we stretch the original trace which lasts for several days and apply random sampling so that the workload we run completes within one hour.
As shown in Figure 6, LMCache consistently outperforms naive vLLM on the real trace on different QPS, reducing both TTFT and ITL, demonstrating its effectiveness in real-world settings. Note that for lower QPS from 2 to 5, LMCache still outperforms naive vLLM by 25%, while when QPS is 6, LMCache has improvement of 49% compared to naive vLLM.
@@ 7.7 Sensitivity Study
Figure 12: Impact of different context lengths.
@@ 8 Real-World Lessons and Experience
8.1 Scaling KV Cache Storage and Offloading
One clear trend in the adoption of LMCache in production systems is that more and more KV cache needs to be stored, definitely beyond the number that can be stored in GPU memory. Early deployments kept all KV caches in GPU memory, but long contexts and multi-user workloads may quickly exhaust GPU memory. Thus, many companies now need to offload KV data to CPU RAM, CPU memory pooling, or even disks when needed. This prevents evicting useful cache entries — without offloading, a follow-up query would trigger full prefill of the same input context, often long contexts, which greatly prolongs TTFT.
Interestingly, some teams, including company R, have experimented with remote storage devices. A remote data fetch is orders of magnitude slower than GPU memory access, yet real-world tests have shown it can still yield both speedups and cost savings. Retrieving previously computed KV cache over the network can be faster than recomputing them on loaded GPUs, with LMCache’s KV cache loading optimizations (§4). While the overhead of loading KV cache from remote is large, it can be pipelined with inference computations, either prefilling requests that do not have cache-hit tokens or decoding computations. These experiences show that even slow and cheap storage can be leveraged to efficiently store and load KV caches.
Another technique used in production is the combination of CPU offloading and PD disaggregation. At company T, on the prefiller side, the KV cache is not only sent to the decoder side, but is also offloaded to the CPU memory of prefiller instance to achieve both the benefit of CPU offloading and PD disaggregation.
8.2 Emerging Use Cases and Surprising Outcomes
While chatbots were one of the first obvious beneficial usage scenarios of KV cache reuse, new use cases are rapidly emerging. Recently, large-scale recommendation systems from company L and company M have also adopted LLMs as the recommender model and becomes an important use case of KV cache reuse. In these systems, LLMs are used as the embedding models to prefill the contexts, without generating tokens, and are frequently reused by different queries of the same user. Furthermore, the contexts are often very long. Caching those KV cache can directly eliminate the expensive prefill phase, thus is a perfect use case of LMCache.
Although multi-round chat has long been a canonical use case for KV cache reuse, real-world deployments have revealed several important insights. Many systems initially adopted sliding context windows to fit more tokens into limited GPU memory during inference. However, this approach often degrades generation quality by discarding useful history and reduces prefix cache hit ratio, since truncated contexts are no longer prefixes. Deployment experiences at Company T and F demonstrate that keeping the full conversation history, while offloading KV caches to larger memory tiers, is a more effective practice. By avoiding truncation, the system achieves significantly higher cache hit rates on shared prefixes, leading to less GPU memory usage per request. Moreover, in production environments, companies have observed that prefix cache hit rate is far higher than they expected.
The experience from Company B’s adoption shows that industry users are accepting lossy optimizations, such as KV cache compression, for better system performance. Especially, in open-ended chatbot applications, where there is no single “correct” answer, using compressed KV cache (e.g., quantized KV cache) may not harm the user experience but can dramatically reduce memory usage and I/O time. Even in finance companies, teams found that slight differences from using a compressed KV cache are tolerable if the overall system throughput improves.
8.3 Integration and Deployment Challenges
Deploying a caching layer in production LLM inference systems has its own practical challenges and lessons. One key lesson we found in the adoption of LMCache is that many companies preferred containerized environments, such as Docker images, when deploying the caching layer into production, as they often do not go into the source code of LMCache.
Another important aspect in the caching layer for LLM inference is fault tolerance and transparency. Because the caching layer sits beneath the inference engines and is supposed to be invisible to end users, robustness is important. Many companies stressed that if the cache system fails or underperforms, it should not bring down the model service. For instance, LMCache incorporates fault-tolerant KV cache retrieval: if errors occur during retrieval, LMCache reports which portions of the KV cache were successfully read, ensuring the inference engine does not break due to the unsuccessful read of KV cache and the correctness of generation is preserved.
We also observed a growing demand for a clean separation between the model serving engine and the caching engine. Many storage companies, such as company I, R, C, and W, have explicitly asked for decoupling the KV cache management and the core inference code, since they want to touch as minimum core inference logic as possible. Although LMCache does not solve this for now, the trend is clear: an LLM inference engine is pairing with a variety of external KV cache management services.
At last, we found that ensuring compatibility between LMCache and emerging inference frameworks was easier than expected. LMCache was originally designed and developed on top of vLLM. However, integrating LMCache with other frameworks, such as SGLang, has also been straightforward, as long as they adopt the paged attention mechanism. Looking ahead, we expect that LMCache will be equally easy to integrate with other inference frameworks based on paged attention.
8.4 Adoption, Lessons Learned, and Future Outlook
Our journey began as an academic prototype in the summer of 2024. At that time, KV caching was a niche concept mostly discussed in a few research papers. But by spring 2025, the landscape had changed — KV cache became a key technique in industry for improving the efficiency in LLM inference. The adoption of LMCache truly exploded in early 2025, with numerous companies starting to collaborate and use.
Interestingly, although LMCache has initially been an academic prototype, we found that LMCache has much more industry users than academia ones. One potential reason is that there is lack of an easy set of programmable interfaces. As a concrete example, a rich set of research literatures explore dropping tokens from KV cache for compressing it into smaller tensors. However, it is non-trivial to implement token dropping in LMCache, since LMCache assumes the number of tokens that is input to and output of LMCache should be the same. Furthermore, any intermediate states in the attention module are hidden from the LMCache’s side, so research works that need to access or modify the attention states are hard to be implemented with LMCache.
Finally, we learned that programming languages for high performance are not always the best fit for building LLM inference infrastructure. Early on, some organizations (e.g.,, company R and I) assumed that KV cache management must be implemented in Rust or C++ to achieve high efficiency. In contrast, we developed LMCache in Python, which turned out to be a more effective choice. Python not only simplifies integration with existing inference frameworks but also lowers the barrier for community contributions, enabling rapid feature development and iterations. That said, CUDA-based implementations are still essential for certain high-performance data loading modules within LMCache.

### v2

@@ 5.2 Compute-I/O Overlapping
LMCache employs multiple optimizations aiming for overlapping LLM inference computations with I/O to maximize GPU utilization.
@@
Asynchronous compute & prefetch:  In many scenarios, there is a time gap between when the inference scheduler admits a query and when the query’s KV cache is actually needed for inference. For example, if a query (with cache hit) arrives when the inference engine is processing other queries, the arriving query has to wait in queue. LMCache exploits this idle interval to prefetch the queued queries’ KV cache from slower storage tiers into faster ones (e.g., from remote disk to local CPU memory or GPU memory). As a result, when the actual inference computation starts, the required KV cache can be loaded or used directly from faster-tier storage, significantly reducing loading delay. LMCache allows users to configure the target tier for prefetching based on their own needs in latency SLO and resource constraints.
@@ 5.3 Minimum Data Copy
Zero-Copy Operations:  When transferring KV cache to multiple devices simultaneously, LMCache minimizes data duplication through a reference counter. Specifically, when KV cache is written to multiple destinations—such as writing from local CPU memory to local disk and remote object storage at the same time—LMCache increments a reference counter on the shared data for each transfer instead of creating new copies. Each completed read or write decrements the counter, and once the count reaches zero, the data is released. This design ensures that data is shared across concurrent read and write operations without unnecessary replication, thus reducing memory pressure and improving efficiency. This technique is simlar to PCB counter in operating systems Strecker1978VAX.
@@
get_num_new_matched_tokens(query) $\rightarrow$ Optional[matched_tokens] Returns the number of cache-hit tokens found in LMCache ’s backend. Returns None if the LMCache decides to let vLLM process other requests first and put this request back to waiting queue.
@@
Note that if a query attempts to allocate pages beyond the current pointer, the allocation must stall until the current pointer moves right enough to cover the required pages. Thus, a key trade-off in this design is that: the number of duplicated pages—i.e.,, the region defined by end pointer – start pointer –between GPU and CPU memory. A smaller duplication window reduces the duplication ratio but increases the likelihood of allocation stalls. For instance, if only one page is duplicated and an inference query requires three pages, the query must wait until the current and end pointers advance by two additional pages. On the other hand, if three pages are duplicated, the same query can proceed immediately without stalling, though at the cost of higher duplication ratio. Though not supported, the same dynamic offloading strategies can also be extended to other storage tiers beyong CPU and GPU.
@@ 6 Standardized Interface for Connecting the KV Caching Layer and Inference Engine
To address this challenge, LMCache introduces a standardized KV cache connector interface that decouples KV cache management from the inference engine backend. This design ensures that LMCache remains compatible regardless of how the upstream inference engine evolves.
We note that the design of this API is initiated by LMCache team, but the implementation and the maintainance of this API are the collaborative effort of both LMCache team and vLLM team.
Design objectives:  The key design objectives are:
  • $\bullet$
    Maximum flexibility: it enables as much KV cache operations as possible.
  • $\bullet$
    vLLM-native: it aligns with the design direction of vLLM, including strict scheduler — worker separation, prefix caching as the first-class citizen, and piece-wise CUDA graph, where vLLM only captures CUDA graph for non-attention operations.
  • $\bullet$
    Friendly to out-of-tree connector: it allow intergrating with out-of-tree connectors without vLLM-side code motification.
  • $\bullet$
    Minimum API-level overhead: it does not introduce overheads (e.g., inter-process communication) at API level.
To be compatible with the scheduler — model runner separation philosophy in vLLM, the connector API contains two sets of interfaces: 1) the scheduler, where the extra cache-hit tokens from the connector are treated as normal prefix-cached tokens in vLLM and directly influences scheduling decisions and are changed by LMCache (i.e., if there is cache hit in LMCache, the number of tokens that need to be newly prefilled changes); and 2) the model runner, where we add hooks before and after model execution, and also before and after attention computation, to enable both bulky KV cache offloading and layer-wise KV cache offloading.
@@
Putting it together, when a query comes in, the scheduler first calls get_num_new_matched_tokens which queries LMCache to see cache hit tokens in the backend. The function can return None if LMCache decides to let vLLM put the current request back to waiting queue and process other requests first, overlapping this request’s I/O with other requests’ computation. Then the update_state_after_alloc function decides whether each page in vLLM needs to be loaded from external storage backend based on matched tokens information from LMCache. If the cache hit tokens are greater than zero, build_connector_meta function is called to prepare necessary metadata to load or store KV cache from storage devices.
@@
Impact:  This API is out for over six months in vLLM. During these six months, we see open-source adoptions, including NVIDIA dynamo project, llm-d project from RedHat, AIBrix project from ByteDance, and vLLM production stack project. We also see multiple proprietary connectors from different companies that use the KV connector API.
Internal APIs Description
batched_admit/batched_evict(hashes, inst_id, device) Send the KV admission/eviction messages from an LMCache instance to the controller manager.
batched_p2p_lookup(hashes) $\rightarrow$ list[inst_id, device, hit_chunks] Lookup peer KV cache existence from an LMCache instance based on the the given hashes.
External APIs Description
lookup(tokens) $\rightarrow$ list[inst_id, device, hit_tokens] Lookup the global KV cache existence of the given tokens.
move((src_inst_id, src_device), (dst_inst_id, dst_device), tokens) Moves the KV cache of the given tokens from source location (src_inst_id, src_device) to destination location (dst_inst_id, dst_device).
clear(tokens, inst_id, device) Clears the KV cache for corresponding tokens from the storage device device in instance inst_id.
pin/unpin(tokens, instance, storage_device) Pins/unpins the KV cache for corresponding tokens at location (inst_id, device).
compress/decompress(tokens, instance, device, method) Compresses/decompresses the KV cache for the corresponding tokens at location (inst_id, device) with a specified compression/decompression method.
Table 3: APIs in LMCache Controller.
@@ 7 Controller Interfaces
LMCache operates as a distributed caching system built around a centralized KV cache controller responsible for global metadata management, cache manipulation, and request routing. To support these functionalities, LMCache provides two categories of APIs: (1) external APIs, which are directly accessible to users or system operators, and (2) internal APIs, which are used by individual LMCache instances.
Mechanically, the KV cache controller consists of two layers: a centralized controller manager and per-instance workers. The controller manager runs as a standalone process and serves as a global coordination point, while per-instance workers collocate with each peer LMCache instance and handle local operations or issue global requests to the manager. External API calls are handled by the centralized manager, which, if necessary, dispatches the appropriate operations to each worker. Per-instance workers can also proactively interact with the centrailzed manager via internal APIs for metadata update or lookup.
The KV cache controller underpins a series of advanced optimizations, including cross-node KV cache sharing, cache-aware request routing, and dynamic KV cache migration. The remainder of this section demonstrates how these optimizations could leverage the controller interfaces through concrete examples.
KV cache–aware routing:  In this case, higher-level routers aim to direct requests to the instance with the highest expected cache hit rate. Each LMCache instance reports its cache admission and eviction decisions to the controller manager via the batched_admit and batched_evict interfaces. The controller manager aggregates these updates and maintains a global in-memory view of the KV cache state across all instances. When the router calls lookup(tokens), the controller consults its in-memory global KV cache states and returns a list of (instance_id, storage_device, hit_tokens), indicating where and how many of the requested tokens are currently cached.
KV cache migration:  When an instance holding KV cache is about to be scaled down or load balancing is required, the KV cache may need to be migrated to another instance. The controller manager hadles such operations through move((src_inst_id, src_deivce), (dst_inst_id, dst_deivce), tokens) API call by dispatching the request to the source instance. The source instance will try to establish connection to the destination instance if one does not exist, and transfers the specified KV cache from the source storage device src_device to the destination location indicated by (dst_inst_id, dst_deivce).
P2P KV cache sharing:  LMCache supports peer-to-peer KV cache sharing, allowing an instance to fetch KV cache from another peer when a local cache miss occurs. Upon a cache miss, the instance’s local worker can query the centralized controller manager via batched_p2p_lookup. The manager will return a list of (inst_id, device, hit_chunks), representing the number of hit chunks and the location that hold these chunks. The instance can then choose to, for example, load KV cache from a peer with maximum hit_chunks.
@@
KV cache clearance:  Applications may clear cache when switching models or reclaiming memory. Upon recieving a clear(tokens, inst_id, location) call, the controller manager dispatches the operation to the corresponding instance identified by inst_id. The instance’s worker then removes the KV cache associated with tokens stored from a specific stroage device device.
@@
Some APIs are not covered in the above applications, such as compress/decompress(tokens, inst_id, device, compression_method) which compresses/decompresses KV cache stored in location (inst_id, device) with a specified compression_method and pin/unpin(tokens, inst_id, device) which can pin/unpin the specifed KV cache at a certain location (inst_id, device). Users can freely call these APIs to explicitly manage the KV cache in their own applications based on their needs.
@@ 8 Evaluation
[…]
@@ 8.1 Setup
Scenario Acronym Single-node / Multi-node Network Medium Real-world Examples
CPU Offload Single-node N.A. Single-node CPU Offloading
Central Storage Single-node Ethernet Centralized Storage Server
PD Single-node NVLink PD Disaggregation
@@
Models:  We compare LMCache against baseline solutions on popular open source models adopted by industry: meta-llama/Llama-3.1-8B-Instruct, Sao10K-L3-8B, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-Coder-32B-Instruct, Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8, Qwen/Qwen2.5-72B-Instruct.
Datasets:  LMCache is evaluated on several datasets, including emulated multi-round question answering, long context question answering from LongBench bai2024longbenchbilingualmultitaskbenchmark, and random dataset from vLLM official benchmarking script vllm.
@@
Baselines:  We compare LMCache v0.3.6 with several baselines, including:
  • $\bullet$
    Basic vLLM: vLLM v0.10.2 which enables prefix caching by default, but only keeps KV cache inside GPU memory, so only a small portion of it can be kept;
  • $\bullet$
    Basic vLLM CPU Offloading: vLLM v0.11.0 with its own implementation of CPU offloading;
  • $\bullet$
    Commercial offerings #1 and #2: provides dedicated endpoint service that reserves GPUs for users to run a user-defined model. We ran these baselines accessed on September 10th.
@@ 8.2 Single-node CPU Offloading
Understanding LMCache’s gains:  LMCache outperforms baselines for several reasons. Compared with basic vLLM, which caches KV data only in GPU memory, LMCache leverages CPU offloading. Since CPU memory can hold far more KV cache than GPU memory, LMCache achieves significantly higher cache hit ratios. Compared to basic vLLM CPU offloading, which employs per-layer and per-16-token transmission and is unable to fully utilize the loading bandwidth, LMCache has more efficient implementation of the transmission module as it loads KV cache at chunk level with high-performance data loading CUDA kernels. Our comparison with closed-source commercial alternatives is conducted in a black-box manner since their internal implementations are not publicly available. From the end-to-end results, we hypothesize that Commercial Option #1 lacks a KV cache offloading mechanism to secondary storage. In contrast, Commercial Option #2 likely supports KV cache offloading to secondary storage, yet its performance is still worse than LMCache.
8.3 Real-trace Driven Evaluation
[…]
We evaluate LMCache on a real trace drawn from company F and company G’s distribution of input and output tokens. Since we do not have access to company $F$’s proprietary models, we run the trace using five different models. To make the experiment tractable, we stretch the original trace which lasts for several days so that the workload we run completes within one hour. LMCache is set to use 500 GB of CPU DRAM at maximum, and we compare it with latest basic vLLM with GPU prefix caching.
[…]
@@ 8.4 Centralized Storage Server
Next, we run LMCache for KV cache sharing through a centralized remote server, that is connected to the GPU instance with a bandwidth of 15 Gbps, following the setup of central storage in Table 4. For this experiment, we evaluate using the TriviaQA dataset from LongBench bai2024longbenchbilingualmultitaskbenchmark, a widely adopted benchmark for long-context evaluation. We follow the official vLLM benchmarking scripts vllm, which generate inference queries according to a Poisson distribution at a specified QPS.
[…]
We note, however, that loading KV cache from the remote backend introduces greater latency than loading from CPU memory, since the remote backend has a much lower bandwidth. As a result, the loading delay may even surpass the prefill delay, particularly when the input context is short or model is small, as prefilling is too fast in such cases—a scenario that we will demonstrate later in 8.7. Thus, adaptive decisions between KV cache loading and prefilling need to be made when KV cache resides in a remote storage server.
@@ 8.5 PD disaggregation
[…]
Figure 14: Compared to vLLM’s native PD disaggregation, LMCache achieves much smaller transmission latency, thus reducing end-to-end delay.
@@ 8.7 Sensitivity Study
Figure 15: At network bandwidth of 32Gbps, LMCache’s KV cache offloading only outperforms basic vLLM’s prefill when input length is more than 256K tokens. At network bandwidth of 64 or 128Gbps, LMCache’s KV cache offloading is better than prefill across all input lengths.
@@ 8.8 SGLang Results
Figure 16: On Qwen3-32B model, LMCache’s CPU offloading achieves comparable performance as SGLang’s native CPU offloading.
@@ 9 Real-World Lessons and Experience
Loading from remote storage is faster than prefill:  Traditionally, it was believed that loading KV cache from remote storage primarily improves cache hit rates and reduces storage costs by leveraging cheaper storage devices, but at the cost of increased inference latency, since loading data from remote devices was thought to be slower than performing a full prefill. This assumption was largely due to the historically low throughput of remote object stores such as Amazon S3, which offered loading speeds as low as 100 MBps. Recently, however, there has been a major boost in remote storage performance, for example, with Amazon S3 Express, the throughput has increased from 100 MBps to nearly 1 GBps. Our users, such as Company C, have adopted LMCache to load KV caches from their own remote object store, achieving 22–32% lower TTFT compared to full prefill. This insight suggests that remote backends can simultaneously improve cache hit ratios and reduce TTFT.
Context truncation lowers prefix cache hit rates:  Many industry users employ a sliding-window mechanism to handle long-context inputs constrained by limited model context windows or GPU memory. For instance, when input tokens exceed the context window limit, some companies truncate the input to keep only the most recent tokens. However, this approach significantly reduces prefix cache hit ratios, since truncated inputs no longer match the prefixes of previously cached contexts. In practice, using real traces from Company F, we observe that prefix cache hit ratios drop from roughly 85% to 45% when truncating input contexts to keep only the latest tokens. Other studies have also discussed this phenomenon, emphasizing that dynamically adding or removing context tokens should be avoided, as it invalidates prefix KV cache reuse manus_blog.
Containerized code is preferred:  With the growing scale of LLM inference, most production environments now rely on Kubernetes to manage GPU clusters. Consequently, deploying inference engines (e.g., vLLM or SGLang) and LMCache through containerized environments, typically via Docker images, has become the standard practice among industry users. Interestingly, many users rely solely on the official Docker images without going deep into LMCache ’s source code.
Unexpected high cache hit rate in production systems:  Our customers did not expect such a high prefix cache hit ratio, such as 50% hit rate for company G in their production environments, until they deploy LMCache inside their systems. Previously, people thought KV caches could only be reused for fixed system prompts. However, modern applications increasingly exhibit "dynamically reusable contexts", such as conversation histories in coding assistants, chat applications, and retrieval-augmented generation (RAG) pipelines. These emerging patterns have significantly increased overall cache hit ratios in real-world deployments.
Industry vs. academia users:  LMCache was designed as a unified prototype framework in May 2024, where we can put our research works into it to gain more impact. However, we found that at that time, industry wants a highly efficient KV cache offloading solution as the size of KV cache and concurrent users keep growing. Then, our attention has been moved towards improving the performance, stability, and compatibility. Since most companies are less concerned with customizing attention algorithms, we deprioritized designing flexible APIs for integrating specialized attention mechanisms, such as selective token dropping. This makes LMCache less popular in academia, since the academia users often focuses on modifying attention mechanisms for their research prototypes. As next step, LMCache is going to design more flexible APIs such that it is easy to use by both industry and academia.
Flexibility vs. performance for programming language:  Python has always been the de facto language in ML. However, the current industry focus is gradually shifting towards higher efficiency instead of broad compatibility. Specifically, many companies are rewriting ML libraries in high-performance languages such as Rust or C++, or carefully optimizing Python-based systems to hide its runtime overhead while preserving flexibility. Although some alternatives have been writing their ML libraries in Rust, we continue to use Python with carefully designed optimizations. This approach allows us to evolve faster, with more community contributions, and still maintain performance on par with alternatives.
LMCache now a community effort:  One key reason LMCache rapidly evolved from a research prototype to a widely adopted industry framework is the active involvement of community contributors. About a year ago, LMCache supported only local CPU, local disk, and Redis backends integrated with vLLM on NVIDIA GPUs. Today, it supports eight more storage backends (NFS, WEKA, GPU-Direct Storage, Mooncake Store, NIXL, S3, InfiniStore, and Valkey) across four processor types (NVIDIA, AMD, Ascend, and TPU), and two inference engines (vLLM and SGLang). All of these contributions were made by industry partners who actively upstreamed their code to stay aligned with ongoing development and avoid divergence from the latest LMCache updates.